# Telegrapher’s Generative Model via Kac Flows

### v4

@@ 5.2 Multi-Dimensional Kac Process
Figure 2: The distribution of the Kac process $\mathbf{X}_{t}$ starting in a 2D Gaussian mixture, simulated until time $T=10$ with $(a,c)=(1,1)$ (upper row) and $(a,c)=(5,5)$ (lower row).
For generative modeling, we propose to circumvent the missing conservation of mass of the multi-dimensional telegrapher’s equation (5), see Remark 6, by using a multivariate random process which consists of independent 1D Kac processes in each spatial component.
For a random variable $\mathbf{X}_{0}\in\mathbb{R}^{d}$, $d\geq 1$, we consider i.i.d. [^2] We use i.i.d. Kac processes just for simplicity. But our framework also allows for independent processes $K_{i}$ each with their own parameters $(a_{i},c_{i})$. copies $K^{i}(t)$, i=1,,di=1,...,d, of a Kac process (25) starting in $\mathbf{0}$. Then, we call
$\mathbf{K}(t):=(K^{1}(t),...,K^{d}(t))$ (64)
a $d$-dimensional Kac process starting in $\mathbf{0}\in\mathbb{R}^{d}$, and
$\mathbf{X}_{t}:=\mathbf{X}_{0}+\mathbf{K}(t)$
the respective Kac process starting in $\mathbf{X}_{0}$.
The stepwise linear movement of the Kac process $(\mathbf{X}_{t})_{t\geq 0}$ translates into a favorable regularity of its probability flow $(P_{\mathbf{X}_{t}})_{t}$. In particular, it is absolutely continuous in the Wasserstein space $\mathcal{P}_{2}(\mathbb{R}^{d})$ and there is no local explosion of the velocity field as in (12).
Proposition 10.
@@
    $\|v_{t}\|_{L_{2}(\mu_{t})}^{2}\leq dc^{2}\quad\text{ for a.e. }t\in[0,\infty),$ (65)
@@
i) Fix $0\leq s<t$. By construction of $\mathbf{X}_{t}$ via (24), we obtain for $\mathbb{P}$-a.e. $\omega$ that
$\|\mathbf{X}_{t}(\omega)-\mathbf{X}_{s}(\omega)\|^{2}$ $=\|\mathbf{K}(t)(\omega)-\mathbf{K}(s)(\omega)\|^{2}=\sum_{i=1}^{d}|K_{i}(t)(\omega)-K_{i}(s)(\omega)|^{2}$
$=\sum_{i=1}^{d}\big|c\,{\rm B_{\frac{1}{2}}}(\omega)\int_{s}^{t}(-1)^{N(z)(\omega)}\,\mathrm{d}z\big|^{2}$
$\leq d\,c^{2}|t-s|^{2}.$ (66)
@@
$\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|(x,y)\|^{2}\,\mathrm{d}\pi(x,y)=\int_{\Omega\times\Omega}\|(\mathbf{X}_{0}+\mathbf{K}(s),\mathbf{X}_{0}+\mathbf{K}(t))\|^{2}\,\mathrm{d}\mathbb{P}$ (67)
$\leq\int_{\Omega\times\Omega}2\|\mathbf{X}_{0}\|^{2}+2\|\mathbf{K}(s)\|^{2}+2\|\mathbf{X}_{0}\|^{2}+2\|\mathbf{K}(t)\|^{2}\,\mathrm{d}\mathbb{P}<\infty,$ (68)
@@
$W_{2}^{2}(\mu_{s},\mu_{t})\leq\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|x-y\|^{2}\,\mathrm{d}\pi(x,y)=\int_{\Omega\times\Omega}\|\mathbf{X}_{s}-\mathbf{X}_{t}\|^{2}\,\mathrm{d}\mathbb{P}\leq d\,c^{2}|t-s|^{2}.$ (69)
@@
ii) By definition (1) and Part i), we see immediately that $(\mu_{t})_{t\geq 0}$ is an absolutely continuous curve. Then it follows from [3, Theorem 8.3.1] that there is a (optimal) velocity field in the continuity equation fulfilling (65).
@@
By the next theorem the convergence result from Theorem 8 stated for the insertion method (57) holds also true for our componentwise Kac process (64).
Theorem 11.
Let $f_{0}$ be a probability density of a random variable $\mathbf{X}_{0}=(X_{0}^{1},...,X_{0}^{d})$ with independent components such that $f_{0}(x)=f_{0}(x^{1},\ldots,x^{d})=\prod_{i=1}^{d}f_{0}^{i}(x^{i})$ and $f_{0}^{i}\in H^{2}(\mathbb{R})$, $i=1,\ldots,d$. Let $\mathbf{X}(t)$ be a $d$-dimensional Kac process starting in $\mathbf{X}_{0}$ with parameters $a>0$ and $c\geq c_{0}>0$ and with probability density $u_{t}$. Further, let $\sigma^{2}:=\frac{c^{2}}{a}$ be fixed and $\mathbf{X}_{0}+\sigma\mathbf{W}(t)$ be a $d$-dimensional Wiener process starting in $\mathbf{X}_{0}$ with probability density $h_{t}$. Assume that $\mathbf{X}_{0}$ is independent of $\mathbf{K}(t)$ and $\sigma\mathbf{W}(t)$, respectively. Then it holds
$\|h(t,\cdot)-u(t,\cdot)\|_{L_{2}(\mathbb{R}^{d})}\leq C_{d}\left(\frac{1}{at^{1+\frac{d-1}{4}}}+\frac{1}{ca^{\frac{1}{2}}t^{\frac{1}{2}+\frac{d-1}{4}}}\right)+R_{d}(a,c,t),$ (70)
where the constant $C_{d}$ depends on $c_{0}$ and $\sigma$, and the function $R_{d}(a,c,t)$ satisfies $\lim_{a,c\uparrow\infty}R_{d}(a,c,t)=0$ for any $t$, and $R_{d}(a,c,\cdot)\in\mathcal{O}(t^{-(\frac{1}{2}+\frac{d}{4})})$ for any $a,c$. Hence, the density of our multi-dimensional Kac process $L_{2}$-converges to the density of the Wiener process with rate $\mathcal{O}(t^{-(\frac{1}{2}+\frac{d-1}{4})})$.
@@
First, by assumption, $(X_{0}^{1},...,X_{0}^{d},K^{1}(t),...,K^{d}(t))$ is (mutually) independent. Hence, $\mathbf{X}_{t}=(X_{0}^{1}+K^{1}(t),...,X_{0}^{d}+K^{d}(t))$ has independent components $X_{0}^{i}+K^{i}(t)$ that are one-dimensional Kac processes starting in $X_{0}^{i}\sim f_{0}^{i}$. By Theorem 1, they admit a probability density $u^{i}$ solving the 1D telegrapher’s equation (5) with initial functions $f_{0}^{i}\in H^{2}(\mathbb{R})$, and hence, $\mathbf{X}_{t}$ admits the probability density $u(t,x)=\prod_{i=1}^{d}u^{i}(t,x^{i})$ by the independence of the components.
Analogously, we conclude that $\mathbf{X}_{0}+\sigma\mathbf{W}(t)$ has the density $h(t,x)=\prod_{i=1}^{d}h^{i}(t,x^{i})$, where the marginal densities $h^{i}$ solve the 1D heat equation (61) with initial functions $f_{0}^{i}\in H^{2}(\mathbb{R})$. Hence, by Theorem 8, the bound (63) holds componentwise with some $C_{1}=C_{1}(c_{0})>0$, i.e.
$\|h^{i}(t,\cdot)-u^{i}(t,\cdot)\|_{L_{2}(\mathbb{R}^{1})}\leq C_{1}\left(\frac{1}{at}+\frac{1}{c(ta)^{\frac{1}{2}}}\right)\quad\text{for all}\quad i=1,...,d.$ (71)
@@
$\prod_{i=1}^{d+1}a_{i}-\prod_{i=1}^{d+1}b_{i}$ $=\big(\prod_{i=1}^{d}a_{i}\big)(a_{d+1}-b_{d+1})+b_{d+1}(\prod_{i=1}^{d}a_{i}-\prod_{i=1}^{d}b_{i}),$ (72)
@@
$\|h(t,\cdot)-u(t,\cdot)\|_{L_{2}(\mathbb{R}^{d+1})}$ $=\Big(\int_{\mathbb{R}^{d+1}}|\prod_{i=1}^{d+1}h^{i}(t,x^{i})-\prod_{i=1}^{d+1}u^{i}(t,x^{i})|^{2}\,\mathrm{d}x\Big)^{\frac{1}{2}}$ (73)
$\leq\Big(\int_{\mathbb{R}^{d+1}}|\big(\prod_{i=1}^{d}h^{i}(t,x^{i})\big)(h^{d+1}(t,x^{d+1})-u^{d+1}(t,x^{d+1}))|^{2}\,\mathrm{d}x\Big)^{\frac{1}{2}}$ (74)
$\quad+\Big(\int_{\mathbb{R}^{d+1}}|u^{d+1}(t,x^{d+1})\big(\prod_{i=1}^{d}h^{i}(t,x^{i})-\prod_{i=1}^{d}u^{i}(t,x^{i})\big)|^{2}\,\mathrm{d}x\Big)^{\frac{1}{2}}$ (75)
$=\Big(\prod_{i=1}^{d}\|h^{i}(t,x^{i})\|_{L_{2}(\mathbb{R}^{1})}\Big)\|h^{d+1}(t,x^{d+1})-u^{d+1}(t,x^{d+1})\|_{L_{2}(\mathbb{R}^{1})}$ (76)
$\quad+\|u^{d+1}(t,x^{d+1})\|_{L_{2}(\mathbb{R}^{1})}\Big(\int_{\mathbb{R}^{d}}|\big(\prod_{i=1}^{d}h^{i}(t,x^{i})-\prod_{i=1}^{d}u^{i}(t,x^{i})\big)|^{2}\,\mathrm{d}x\Big)^{\frac{1}{2}}.$ (77)
@@
$\frac{C_{1}}{(4\pi\sigma^{2})^{\frac{d}{4}}}\left(\frac{1}{at^{1+\frac{d}{4}}}+\frac{1}{ca^{\frac{1}{2}}t^{\frac{1}{2}+\frac{d}{4}}}\right).$ (78)
@@
$\quad\Big(\|u^{d+1}(t,x^{d+1})-h^{d+1}(t,x^{d+1})\|_{L_{2}(\mathbb{R}^{1})}+\|h^{d+1}(t,x^{d+1})\|_{L_{2}(\mathbb{R}^{1})}\Big)$ (79)
$\quad\cdot\Big(C_{d}\Big(\frac{1}{at^{1+\frac{d-1}{4}}}+\frac{1}{ca^{\frac{1}{2}}t^{\frac{1}{2}+\frac{d-1}{4}}}\Big)+R_{d}(a,c,t)\Big)$ (80)
$\leq\Big(C_{1}\Big(\frac{1}{at}+\frac{1}{c(ta)^{\frac{1}{2}}}\Big)+\frac{1}{(4\pi\sigma^{2}t)^{\frac{1}{4}}}\Big)$ (81)
$\quad\cdot\Big(C_{d}\Big(\frac{1}{at^{1+\frac{d-1}{4}}}+\frac{1}{ca^{\frac{1}{2}}t^{\frac{1}{2}+\frac{d-1}{4}}}\Big)+R_{d}(a,c,t)\Big).$ (82)
@@
The above assumption on the independence of the initial components $(X_{0}^{1},...,X_{0}^{d})$ was made for simplicity reasons. The convergence result for $t\to\infty$ appears to hold true more generally and we forgo the technical details. Furthermore, weak convergence results concerning the paths of a rescaled Kac process towards a Brownian motion can be found in [14, 38].
Figure 3: The distribution of $\mathbf{X}_{t}$ starting in a 2D Gaussian mixture, simulated until time $T=3$ with ’large’ parameters $(a,c)=(100,10)$. As described in Theorem 11, we see a ’diffusion’ like process with approximate variance $\sigma^{2}t=\frac{c^{2}}{a}t=t$.

### v5

@@ 5.2 Multi-Dimensional Kac Process
Figure 2: The distribution of the Kac process $\mathbf{X}_{t}$ starting in a 2D Gaussian mixture, simulated until time $T=10$ with $(a,c)=(1,1)$ (upper row) and $(a,c)=(5,5)$ (lower row). We observe a so-called ballistic-to-diffusive crossover of the Kac process, see e.g. [31].
For generative modeling, we propose to circumvent the lack of mass conservation of the multi-dimensional telegrapher’s equation (5.1), see Remark 5.2, by using a multivariate random process which consists of independent 1D Kac processes in each spatial component – in complete analogy to multi-dimensional Brownian motions.
More precisely, we consider two types of componentwise processes, commonly referred to as variance-exploding (VE) and variance-preserving (VP). For both, we prove Lipschitz regularity in the Wasserstein-2 space in the Propositions 5.6 and A.1. Typically, in the VE case, the latent distribution is obtained by sampling the marginal distribution of the process in the long-time limit $t\to\infty$. Hence, mostly relevant for the VE case, we prove the convergence Theorem 5.7 based on Theorem 5.4.
For a random variable $\mathbf{X}_{0}\in\mathbb{R}^{d}$, $d\geq 1$, we consider i.i.d. [^2] We use i.i.d. Kac processes just for simplicity. But our framework also allows for independent processes $K^{i}$ each with their own parameters $(a_{i},c_{i})$. copies $K^{i}(t)$, i=1,,di=1,...,d, of a Kac process (4.6) starting in $\mathbf{0}$. Then, we call
$\mathbf{K}(t):=(K^{1}(t),...,K^{d}(t))$ (5.10)
a $d$-dimensional Kac process starting in $\mathbf{0}\in\mathbb{R}^{d}$. Corresponding to the VE case, we consider the respective Kac process starting in $\mathbf{X}_{0}$, or VE Kac process, given as
$\mathbf{X}_{t}:=\mathbf{X}_{0}+\mathbf{K}(t),\quad t\in[0,\infty).$ (5.11)
For the VP case, we define the mean‐reverting Kac process, or VP Kac process, by
$\mathbf{M}_{t}\coloneqq(1-t)\,\mathbf{X}_{0}+\mathbf{K}(t),\quad t\in[0,1].$ (5.12)
The stepwise linear movement of the Kac process $(\mathbf{K}(t))_{t\geq 0}$ translates into a favorable regularity of both probability flows $(P_{\mathbf{X}_{t}})_{t}$ and $(P_{\mathbf{M}_{t}})_{t}$. In particular, they are absolutely continuous in the Wasserstein space $\mathcal{P}_{2}(\mathbb{R}^{d})$ and there is no local explosion of the velocity field as in (3.7). We first state the result for the VE Kac process (5.11).
Proposition 5.6.
@@
    $\|v_{t}\|_{L_{2}(\mu_{t})}^{2}\leq dc^{2}\quad\text{ for a.e. }t\in[0,\infty),$ (5.13)
@@
i) Fix $0\leq s<t$. By construction of $\mathbf{X}_{t}$ via (4.5), we obtain for $\mathbb{P}$-a.e. $\omega$ that
$\|\mathbf{X}_{t}(\omega)-\mathbf{X}_{s}(\omega)\|^{2}$ $=\|\mathbf{K}(t)(\omega)-\mathbf{K}(s)(\omega)\|^{2}=\sum_{i=1}^{d}|K^{i}(t)(\omega)-K^{i}(s)(\omega)|^{2}$
$=\sum_{i=1}^{d}\big|c\,{\rm B}_{\frac{1}{2}}^{i}(\omega)\int_{s}^{t}(-1)^{N^{i}(z)(\omega)}\,\mathrm{d}z\big|^{2}$
$\leq d\,c^{2}|t-s|^{2}.$ (5.14)
@@
$\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|(x,y)\|^{2}\,\mathrm{d}\pi(x,y)=\int_{\Omega}\|(\mathbf{X}_{0}+\mathbf{K}(s),\mathbf{X}_{0}+\mathbf{K}(t))\|^{2}\,\mathrm{d}\mathbb{P}$ (5.15)
$\leq\int_{\Omega}2\|\mathbf{X}_{0}\|^{2}+2\|\mathbf{K}(s)\|^{2}+2\|\mathbf{X}_{0}\|^{2}+2\|\mathbf{K}(t)\|^{2}\,\mathrm{d}\mathbb{P}<\infty,$ (5.16)
@@
$W_{2}^{2}(\mu_{s},\mu_{t})\leq\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|x-y\|^{2}\,\mathrm{d}\pi(x,y)=\int_{\Omega}\|\mathbf{X}_{s}-\mathbf{X}_{t}\|^{2}\,\mathrm{d}\mathbb{P}\leq d\,c^{2}|t-s|^{2}.$ (5.17)
@@
ii) By definition (2.1) and Part i), we see immediately that $(\mu_{t})_{t\geq 0}$ is an absolutely continuous curve. Then it follows from [3, Theorem 8.3.1] that there is a (optimal) velocity field in the continuity equation fulfilling (5.13).
@@
Similar to the VE Kac process (5.11), the VP Kac process (5.12) also admits a Lipschitz property in $\mathcal{P}_{2}$ and hence a bounded velocity norm $\|v_{t}\|_{L_{2}(\mu_{t})}$. The statement and proof is given in Proposition A.1 in the appendix.
By the next theorem the convergence result from Theorem 5.4 stated for the insertion method (5.3) holds also true for the VE Kac process (5.11).
Theorem 5.7.
Let $f_{0}$ be a probability density of a random variable $\mathbf{X}_{0}=(X_{0}^{1},...,X_{0}^{d})$ with independent components such that $f_{0}(x)=f_{0}(x^{1},\ldots,x^{d})=\prod_{i=1}^{d}f_{0}^{i}(x^{i})$ and $f_{0}^{i}\in H^{2}(\mathbb{R})$, $i=1,\ldots,d$. Let $\mathbf{X}_{t}$ be a $d$-dimensional Kac process starting in $\mathbf{X}_{0}$ with parameters $a>0$ and $c\geq c_{0}>0$ and with probability density $u_{t}$. Further, let $\sigma^{2}:=\frac{c^{2}}{a}$ be fixed and $\mathbf{X}_{0}+\sigma\mathbf{W}(t)$ be a $d$-dimensional Wiener process starting in $\mathbf{X}_{0}$ with probability density $h_{t}$. Assume that $\mathbf{X}_{0}$ is independent of $\mathbf{K}(t)$ and $\sigma\mathbf{W}(t)$, respectively. Then it holds
$\|h(t,\cdot)-u(t,\cdot)\|_{L_{2}(\mathbb{R}^{d})}\leq C_{d}\left(\frac{1}{at^{1+\frac{d-1}{4}}}+\frac{1}{ca^{\frac{1}{2}}t^{\frac{1}{2}+\frac{d-1}{4}}}\right)+R_{d}(a,c,t),$ (5.18)
where the constant $C_{d}$ depends on $c_{0}$ and $\sigma$, and the function $R_{d}(a,c,t)$ satisfies $\lim_{a,c\uparrow\infty}R_{d}(a,c,t)=0$ for any $t$, and $R_{d}(a,c,\cdot)\in\mathcal{O}(t^{-(\frac{1}{2}+\frac{d}{4})})$ for any $a,c$. Hence, the density of the VE Kac process (5.11) $L_{2}$-converges to the density of the Wiener process with rate $\mathcal{O}(t^{-(\frac{1}{2}+\frac{d-1}{4})})$.
@@
First, by assumption, $(X_{0}^{1},...,X_{0}^{d},K^{1}(t),...,K^{d}(t))$ is (mutually) independent. Hence, $\mathbf{X}_{t}=(X_{0}^{1}+K^{1}(t),...,X_{0}^{d}+K^{d}(t))$ has independent components $X_{0}^{i}+K^{i}(t)$ that are one-dimensional Kac processes starting in $X_{0}^{i}\sim f_{0}^{i}$. By Theorem 4.1, they admit a probability density $u^{i}$ solving the 1D telegrapher’s equation (5.1) with initial functions $f_{0}^{i}\in H^{2}(\mathbb{R})$, and hence, $\mathbf{X}_{t}$ admits the probability density $u(t,x)=\prod_{i=1}^{d}u^{i}(t,x^{i})$ by the independence of the components.
Analogously, we conclude that $\mathbf{X}_{0}+\sigma\mathbf{W}(t)$ has the density $h(t,x)=\prod_{i=1}^{d}h^{i}(t,x^{i})$, where the marginal densities $h^{i}$ solve the 1D heat equation (5.7) with initial functions $f_{0}^{i}\in H^{2}(\mathbb{R})$. Hence, by Theorem 5.4 and Remark 5.5, the bound (5.9) holds componentwise with some $C_{1}=C_{1}(c_{0})>0$, i.e.
$\|h^{i}(t,\cdot)-u^{i}(t,\cdot)\|_{L_{2}(\mathbb{R}^{1})}\leq C_{1}\left(\frac{1}{at}+\frac{1}{c(ta)^{\frac{1}{2}}}\right)\quad\text{for all}\quad i=1,...,d.$ (5.19)
@@
$\prod_{i=1}^{d+1}a_{i}-\prod_{i=1}^{d+1}b_{i}$ $=\big(\prod_{i=1}^{d}a_{i}\big)(a_{d+1}-b_{d+1})+b_{d+1}(\prod_{i=1}^{d}a_{i}-\prod_{i=1}^{d}b_{i}),$ (5.20)
@@
$\|h(t,\cdot)-u(t,\cdot)\|_{L_{2}(\mathbb{R}^{d+1})}$ $=\Big(\int_{\mathbb{R}^{d+1}}|\prod_{i=1}^{d+1}h^{i}(t,x^{i})-\prod_{i=1}^{d+1}u^{i}(t,x^{i})|^{2}\,\mathrm{d}x\Big)^{\frac{1}{2}}$ (5.21)
$\leq\Big(\int_{\mathbb{R}^{d+1}}|\big(\prod_{i=1}^{d}h^{i}(t,x^{i})\big)(h^{d+1}(t,x^{d+1})-u^{d+1}(t,x^{d+1}))|^{2}\,\mathrm{d}x\Big)^{\frac{1}{2}}$ (5.22)
$\quad+\Big(\int_{\mathbb{R}^{d+1}}|u^{d+1}(t,x^{d+1})\big(\prod_{i=1}^{d}h^{i}(t,x^{i})-\prod_{i=1}^{d}u^{i}(t,x^{i})\big)|^{2}\,\mathrm{d}x\Big)^{\frac{1}{2}}$ (5.23)
$=\Big(\prod_{i=1}^{d}\|h^{i}(t,x^{i})\|_{L_{2}(\mathbb{R}^{1})}\Big)\|h^{d+1}(t,x^{d+1})-u^{d+1}(t,x^{d+1})\|_{L_{2}(\mathbb{R}^{1})}$ (5.24)
$\quad+\|u^{d+1}(t,x^{d+1})\|_{L_{2}(\mathbb{R}^{1})}\Big(\int_{\mathbb{R}^{d}}|\big(\prod_{i=1}^{d}h^{i}(t,x^{i})-\prod_{i=1}^{d}u^{i}(t,x^{i})\big)|^{2}\,\mathrm{d}x\Big)^{\frac{1}{2}}.$ (5.25)
@@
$\frac{C_{1}}{(4\pi\sigma^{2})^{\frac{d}{4}}}\left(\frac{1}{at^{1+\frac{d}{4}}}+\frac{1}{ca^{\frac{1}{2}}t^{\frac{1}{2}+\frac{d}{4}}}\right).$ (5.26)
@@
$\quad\Big(\|u^{d+1}(t,x^{d+1})-h^{d+1}(t,x^{d+1})\|_{L_{2}(\mathbb{R}^{1})}+\|h^{d+1}(t,x^{d+1})\|_{L_{2}(\mathbb{R}^{1})}\Big)$ (5.27)
$\quad\cdot\Big(C_{d}\Big(\frac{1}{at^{1+\frac{d-1}{4}}}+\frac{1}{ca^{\frac{1}{2}}t^{\frac{1}{2}+\frac{d-1}{4}}}\Big)+R_{d}(a,c,t)\Big)$ (5.28)
$\leq\Big(C_{1}\Big(\frac{1}{at}+\frac{1}{c(ta)^{\frac{1}{2}}}\Big)+\frac{1}{(4\pi\sigma^{2}t)^{\frac{1}{4}}}\Big)$ (5.29)
$\quad\cdot\Big(C_{d}\Big(\frac{1}{at^{1+\frac{d-1}{4}}}+\frac{1}{ca^{\frac{1}{2}}t^{\frac{1}{2}+\frac{d-1}{4}}}\Big)+R_{d}(a,c,t)\Big).$ (5.30)
@@
The above assumption on the independence of the initial components $(X_{0}^{1},...,X_{0}^{d})$ was made for simplicity. The convergence result for $t\to\infty$ appears to hold true more generally and we forgo the technical details. Furthermore, weak convergence results concerning the paths of a rescaled Kac process towards a Brownian motion can be found in [19, 42].
Lastly, we note that the long-time asymptotics of Theorem 5.7 only reasonably apply to the VE Kac process (5.11), since the VP Kac process (5.12) is constructed to reach its terminal distribution $P_{\mathbf{K}(1)}$ in finite time. [^3] But note that Theorem 5.7 also describes the asymptotic behaviour of the Kac process in the limit $a,c\uparrow\infty$.
Figure 3: The distribution of $\mathbf{X}_{t}$ starting in a 2D Gaussian mixture, simulated until time $T=3$ with ’large’ parameters $(a,c)=(100,10)$. As described in Theorem 5.7, we see a ’diffusion’ like process with approximate variance $\sigma^{2}t=\frac{c^{2}}{a}t=t$.